The distance travelled by a body dropped from the top of a tower is proportional to?

Grade:10

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

S = ut + 0.5at*t
Here u is 0 a is g

S = 0.5 gt*t

S is proportional to square of time t

Regards
Khimraj
4 years ago
 
Newton’s Second Equation 
h=ut + 1/2(at2)
for free fall u = 0 a=g and t=n  g is gravity t is time u is initial velocity
therfore h=0+1/2(gn2)
h is directly proportional to n2